OpenStudy (anonymous):

Show that \[\left(\sum_{k=0}^\infty x^k\right)^2=\sum_{k=0}^\infty (k+1)x^{k}\] I encourage anyone to try to find a "closed" form for \[\left(\sum_{k=0}^\infty x^k\right)^n\] so that a sum raised to a power is equal to another sum not raised to a power. \((n\in\mathbb{N})\)

OpenStudy (anonymous):

Hmm, I think ultimately it would involve integration.

OpenStudy (anonymous):

The right side would only converge if \(x<1\), I believe. It's a geometric series.

OpenStudy (anonymous):

Right, let's also assume \(|x|<1\) for convenience!

OpenStudy (anonymous):

At first I had a tedious (sort of combinatoric) approach in mind that involves expanding the polynomial \((1+x+x^2+\cdots)^2\) and I noticed the pattern above, so I was wondering if there's one for any natural power \(n\). We could also do this by recognizing that the LHS is the power series for \(\dfrac{1}{(1-x)^2}\) and integrating/differentiating the RHS to establish equality, which seems to be the direction you're going in.

OpenStudy (anonymous):

\[ \left(\sum_{k=0}^\infty x^k\right)^n = \left(\frac{1}{1-x}\right)^n =(1-x)^{-n} \]

OpenStudy (anonymous):

For \(n=3\) we get \[\left(\sum_{k=0}^\infty x^k\right)^3=\sum_{k=0}^\infty \frac{(k+1)(k+2)}{2}x^k\] and for \(n=4\), \[\left(\sum_{k=0}^\infty x^k\right)^3=\sum_{k=0}^\infty \frac{(k+1)(k+2)(2k+3)}{6}x^k\] with those coefficients representing triangular and tetrahedral numbers, \[\sum_{r=1}^{k+1}r^2=\frac{(k+1)(k+2)}{2}\quad\text{and}\quad\sum_{r=1}^{k+1}=\frac{(k+1)(k+2)(2k+3)}{6}\] so there might not be a "closed" form for any \(n\)...

OpenStudy (anonymous):

Not that I'm aware of, at any rate.

OpenStudy (anonymous):

Oops,\[\sum_{r=1}^{k+1}\color{red}r=\frac{(k+1)(k+2)}{2}\quad\text{and}\quad\sum_{r=1}^{k+1}\color{red}{r^2}=\frac{(k+1)(k+2)(2k+3)}{6}\]

OpenStudy (anonymous):

Hmm, is the binomial theorem valid for negative powers?

OpenStudy (anonymous):

We can say this much I think: \[\left(\sum_{k=0}^\infty x^k\right)^n=\sum_{k=0}^\infty \left(\sum_{r=1}^{k+1}r^{n-2}\right)x^k\]

OpenStudy (anonymous):

I'm taking Maclaurin series for \((1-x)^{-n}\)

OpenStudy (anonymous):

\[ f^{(0)}(0) = 1\\ f^{(1)}(0) = n\\ f^{(2)}(0) = (n+1)n\\ f^{(3)}(0) = (n+2)(n+1)n\\ f^{(k)}(0) = (n+k-1)\ldots (n+1)n = \frac{(n+k-1)!}{(n-1)!}\\ \]

OpenStudy (anonymous):

I think this means: \[ (1-x)^{-n}=\sum_{k=0}^{\infty }\frac{(n+k-1)!}{(n-1)!} \frac{x^k}{k!} \]

OpenStudy (anonymous):

But that would mean: \[ (1-x)^{-n}=\sum_{k=0}^{\infty }\frac{(n+k-1)!}{(n-1)!} \frac{x^k}{k!} = \sum_{k=0}^{\infty }{n+k-1\choose k}x^k \]

OpenStudy (anonymous):

Which would make your equality seem a bit strange.

OpenStudy (anonymous):

Nice, the formula seems to check out in Mathematica

OpenStudy (anonymous):

For \(n=2\), we get \[\sum_{k=0}^\infty \binom {k+1}kx^k=\sum_{k=0}^\infty (k+1)x^k\]